OpenStudy (anonymous):

Multiply and simplify: (9+square root of 3)^2 A. 81+2*square root of 3 B. 84+18*square root of 3 C. 84

OpenStudy (anonymous):

84+18*root 3

OpenStudy (anonymous):

How'd u get tht...do u mind showing ur wrk?

OpenStudy (raspberryjam):

\[(9+\sqrt{3})^2=(9+\sqrt{3})(9+\sqrt{3})\] \[(81+9\sqrt{3}+9\sqrt{3}+3) = (84+18\sqrt{3})\]

OpenStudy (anonymous):

yeah wut that guy did thats correct

OpenStudy (anonymous):

lol k thnks